# Evaluating LLMs at Evaluating Temporal Generalization

### v1

@@ 3.2 Temporal Bias
Table 1: Comprehensive list of models with their release dates and metrics for TBI (multiplied by 1000) across the arXiv, BBC, and Wiki datasets. Negative slope values indicate a trend of increasing BPC, which refers to decreasing performance.
3.3 Temporal Generalization
To quantify model temporal generalization, we computed the BPC for various models on the Wiki and BBC datasets. We established a base BPC using data from the six months prior to each model’s release. Subsequently, we measured the changes in BPC at 3, 6, 9 and 12 months post-release relative to the base BPC. The specific results are detailed in Appendix C
The generalization of BPC in Wikipedia dataset shows that the models exhibit a relatively stable performance on Wikipedia data, which suggests a consistent ability to handle well-structured and formal text types like those found in Wikipedia.
The generalization of BPC in BBC dataset shows that, as opposed to Wikipedia data, providing a more significant differentiation between models. This distinction is crucial as it reveals the models’ ability to generalize to future, unseen data. Notably, although there is an overall trend of increasing BPC for BBC text from 2020 to 2024, the immediate post-cutoff performance of individual models varies. For instance, Baichuan2 series appears to adapt better to new texts, showcasing an improved familiarity with them. On the other hand, models like Zephyr exhibit poorer generalization in this context.
In our analysis across both the Wiki and BBC datasets, we have observed significant correlations between model attributes and performance decay, detailed in Table 2. The data shows a consistent pattern where higher benchmark scores correlate with faster performance decay, indicating a trade-off between capability and flexibility. This suggests that models with higher capabilities and sophistication, as reflected in their task scores, may compromise adaptability, becoming so specialized that they struggle to adjust to new or evolving data scenarios without a noticeable decrease in performance.
Specifially, in the Wiki dataset, larger models, which generally have higher capabilities, tend to decay quicker (with a correlation of 0.504), likely due to overfitting on training data. Conversely, in the BBC dataset, where the information is more dynamic and rapidly changing, the correlation between model size and performance decay is almost negligible, suggesting that the impact of model size on decay is significantly less in environments that continuously present new challenges. Additionally, across both datasets, newer models show larger correlations, hinting that despite their advanced capabilities, they may decay faster, potentially as a consequence of these enhanced capabilities.
These insights emphasize the importance of designing models that not only perform well on benchmarks but also maintain robustness over time and across varying data environments. As machine learning applications continue to grow in complexity and scale, the ability to generalize well temporally will become increasingly critical. This analysis serves as a cornerstone for future research aimed at enhancing the longevity and adaptability of machine learning models in dynamic settings.
Task or model feature Correlation with BPC change rate average (Wiki) Correlation with BPC change rate average (BBC)
Model Size 0.504 -0.044
Release Date(month since 2015) 0.267 0.313
MMLU 0.599 0.384
MMLU:Humanities 0.500 0.317
MMLU:SocialSciences 0.594 0.311
MMLU:STEM 0.393 0.186
MMLU:Other 0.528 0.153
GSM8 0.473 0.188
Ar20c 0.380 -0.071
Hellawasg 0.581 0.248
TrustfulQA 0.200 0.485
Winogrande 0.632 0.204
Longbench 0.404 0.549
Table 2: Combined correlation between task scores and performance decay on Wikipedia and BBC
3.4 Correlation with Existing Benchmarks
@@
HellaSwag Tests an AI’s common sense reasoning by requiring it to complete sentences or narratives that reflect everyday logic, highlighting the model’s ability to predict logical continuations. (Zellers et al., 2019)
MMLU Evaluates an AI’s breadth and depth of understanding across a diverse range of subjects, from science to literature, showcasing its versatility and comprehensive grasp of human language. (Hendrycks et al., 2021)
TruthfulQA Focuses on the model’s capacity to provide honest and accurate responses, particularly to questions where the truth may not be intuitive, testing its commitment to truthfulness and its proficiency in avoiding misinformation. (Lin et al., 2022)
Winogrande Assesses common sense reasoning and natural language understanding through sentence completion tasks, measuring an AI’s ability to apply common sense knowledge in language processing. (Sakaguchi et al., 2019)
GSM8K Challenges AI models with elementary-level math problems, testing their mathematical reasoning capabilities and their understanding and application of basic math concepts. (Cobbe et al., 2021)
Table 3: Overview of AI Benchmarks and Their Evaluation Focus
@@
Table 4 shows the comparative analysis of base and tuned models. Our findings suggest that: 1) Base models exhibit a stronger correlation with established benchmarks, implying that these models, in their fundamental form, possess a robust capacity for modeling language. This observation advocates for conducting initial evaluations using base models to capture their intrinsic language modeling capabilities. 2) The exploration of tuned models reveals potential deeper patterns in model tuning, suggesting that specific tuning strategies may unlock further enhancements in model performance across different tasks and domains.
@@
Table 4: Correlation with Different Benchmarks, Base vs. Chat Models
@@
Table 5 is about a long text understanding benchmark Longbench.
@@
Table 5: Correlation with Longbench values across different lengths
@@
As can be seen in Appendix  D, whether all models are observed or only those of a specific scale does not make a big difference. Comparisons across different model sizes show that the differences are not substantial, indicating consistency in performance correlations across benchmarks.
Further analysis from Table 6 within the MMLU benchmark sections reveals interesting patterns. While we assume models to be generally excel in domains closely aligned with their training data, this correlation is not strong. Specifically, STEM scores do not show strong correlation with STEM-related texts. Moreover, academic papers, even in Social Sciences and Humanities, exhibit stronger correlations than Internet data, which includes forums and news.
@@
3.5 The Drawbacks of Evaluating Language Likelihood
Inaccessibility to API-based Models
A fundamental limitation of using language likelihood as a metric lies in its inapplicability to closed-source models, where key indicators such as perplexity, logits, or losses are inaccessible. This restriction excludes a broad spectrum of proprietary models from evaluation and significantly narrows the scope of comparative analysis. The consequent inability to assess these models hampers our ability to gain a comprehensive understanding of the language technology landscape and the capabilities of different models within it.
Susceptibility to Training Focus
The effectiveness of language likelihood as a metric can be compromised by models that are finely tuned or trained with a particular emphasis on certain formats or domains. In such cases, the likelihood measure might disproportionately favor these specialized areas, skewing the overall evaluation and presenting a distorted view of a model’s general performance.
4 Temporal Generalization in terms of Prognostication Prediction
The above Drawbacks in Sec. 3.5 motivated us to research temporal generalization and bias from the perspective of future prediction. In the context of the model release date, events that occurred before this date are considered ”past events”, while those that took place after are termed ”future events”. In this chapter, we have divided the experiments into two categories based on this criterion: Bias and Temporal Generation.
4.1 Experimental Settings
4.1.1 Data Source
We have sourced all questions from  Good Judgment Open and all binary questions from  Metaculus through web scraping and API downloads. The collected data includes initiation and conclusion dates, domains, questions, descriptions, and answers, which are ordered appropriately for further processing.
4.1.2 Extraction of Predictive Responses
Using the lm-eval-harness  (Gao et al., 2023) tool, we predict model choices based on probability associated with each option. This probability-driven method serves as a reliable metric for evaluating the predictive performance of language models across a variety of topics and question formats.
4.2 Temporal Bias
As can be seen in Table 7, most analyzed models exhibit a negative slope. This is an uncommon scenario in a typical statistical analysis, indicating a general tendency across models to perform better or be more aligned with historical data than recent data. We define the following classification thresholds to discern subtle differences in how models handle temporal information, despite an overall inclination towards nostalgia.
  • Positive TBI (Neophilia Bias): TBI values greater than 0.001 suggest the model has a neophilia bias. This implies the model’s performance is better on more recent data, potentially favoring newer over older training data.
  • Near-Zero TBI (Balanced): TBI values between -0.001 and 0.001 indicate a balanced performance over time. A TBI of 0, in particular, suggests that the model’s performance is evenly distributed around the mean, showing no temporal bias.
  • Negative TBI (Nostalgia Bias): TBI values less than -0.001 indicate a nostalgia bias, suggesting the model may favor historical data, possibly overlooking important recent information.
The choice of ±0.001 as the threshold for these categories is designed to fine-tune the sensitivity of our analysis to subtle temporal biases, ensuring that the classification is both meaningful and practically significant.
model TBI y-axis intercept bias category
OPT-13B -1.37 0.70 Nostalgia
OPT-2.7B -2.85 0.97 Nostalgia
LLaMA-7B -3.71 0.96 Nostalgia
Pythia-12B -3.90 0.95 Nostalgia
Falcon-rw-1B -2.47 0.84 Nostalgia
Baichuan-13B-Chat -2.79 0.98 Nostalgia
Baichuan-7B-Chat -1.17 0.80 Nostalgia
GPT-3.5-turbo-0613 -0.81 0.86 Symmetrical
LLaMA-2-7B -1.66 0.75 Nostalgia
LLaMA-2-13B -0.80 0.73 Symmetrical
LLaMA-2-7B-Chat -3.92 0.93 Nostalgia
Zhongjing-Base 0.08 0.63 Symmetrical
InternLM-Chat-7B -1.76 0.94 Nostalgia
Baichuan2-7B-Base -3.50 0.84 Nostalgia
Baichuan2-7B-Chat -3.71 0.94 Nostalgia
Mistral-7B-v0.1 -0.47 0.75 Symmetrical
Phi-1.5 0.24 0.38 Symmetrical
Baichuan2-13B-Base -3.06 0.72 Nostalgia
Baichuan2-13B-Chat -1.53 0.70 Nostalgia
Colossal-LLaMA-2-7B-Base -3.38 0.91 Nostalgia
Qwen-14B-Chat -2.31 0.93 Nostalgia
Qwen-7B -0.35 0.69 Symmetrical
Qwen-7B-Chat -2.98 0.62 Nostalgia
Skywork-13B-Base -1.46 0.65 Nostalgia
ChatGLM3-6B -4.46 0.92 Nostalgia
Zephyr-7B-beta -2.73 0.99 Nostalgia
Yi-6B -5.18 0.98 Nostalgia
Yi-6B-Chat 0.34 0.71 Symmetrical
Qwen-1.8B -3.59 0.80 Nostalgia
Qwen-1.8B-Chat -4.32 0.96 Nostalgia
GPT-4-1106 -0.08 0.96 Symmetrical
TinyLLaMA-1.1B-Chat-v0.6 -1.78 0.53 Nostalgia
Gemini-pro-1213 -0.43 0.76 Symmetrical
Phi-2 -1.54 0.77 Nostalgia
Table 7: Predict performance of models: TBI(slope) and y-axis intercept are linear fit parameters, bias category classifies the nature of the bias based on TBI
Table 8 shows an analysis covering the period from 2015 to 2024, with the year 2019 serving as a pivotal midpoint, a distinct pattern emerges from the performance data of various language models: all models reached their peak accuracy before 2020. This observation points to a notable trend toward ”nostalgia,” where models demonstrate optimal performance on data from the earlier part of the timeline analyzed. This tendency highlights the models’ alignment with the informational characteristics prevalent prior to the latter half of the observed period, suggesting their training data may disproportionately reflect earlier times.
Peak Period Model Peak Time of Accuracy Cutoff Time Interpolation
Post-2019 TinyLLaMA1.1B 2019-07 2023-11 52
2017 to 2018 GPT4-1106 2018-07 2023-06 59
LLaMA2-13B 2018-07 2023-02 55
Baichuan2-13B-Base 2018-01 2023-06 65
2015 to 2016 Baichuan2-7B-Chat 2016-01 2023-08 91
Colossal-LLaMA-2-7B-Base 2016-01 2023-09 92
LLaMA2-7B 2016-01 2023-02 85
4.3 Temporal Generalization
It is crucial to note that for models, any data post-release represents the future, allowing us to assess their generalization capabilities on upcoming events.
In the analysis of temporal generalization, it is important to acknowledge that the performance of models post-release primarily reflects their ability to predict future events based on past training. As evidenced by the data in Table 9, many models show significant fluctuations in accuracy percentages following their release. This variability suggests that immediately after the cutoff event, most models are essentially guessing, leading to high volatility in performance metrics.
Given the substantial fluctuations observed post-release, the temporal generalization of current models may appear to have limited significance. This is largely because these models exhibit weak capabilities in adapting to new, unseen data, often resulting in what can be described as near-random performance at the cutoff point. Such outcomes highlight the models’ limited understanding of the world and their inability to effectively generalize from past information to future contexts.
However, this should not detract from the potential utility of these measures in the future. As language models continue to evolve and become more sophisticated in their world understanding, the metric of temporal generalization will likely become more relevant and insightful. Future models that better comprehend and adapt to the world will provide more meaningful insights into their generalization capabilities over time, thereby enhancing the value of this assessment.
The current limitations underscore the need for further development in language modeling, particularly in improving the models’ ability to generalize across time. As we progress towards models with a deeper and more accurate understanding of complex world dynamics, temporal generalization measures will play a crucial role in evaluating their long-term effectiveness and adaptability.
Base adjusted ACC Adjusted ACC Change at 3 months (%) Adjusted ACC Change at 6 months (%) Adjusted ACC Change at 9 months (%) adjusted ACC Change at 12 months (%) Mean (%)
Model
OPT-2.7B 0.86 0.11 -13.63 -23.41 -18.98 -13.98
LLaMA-7B 0.66 -0.31 -20.59 5.09 9.89 -1.48
Pythia-12B 0.77 -27.68 -24.85 -6.22 -3.68 -15.61
Falcon-rw-1B 0.69 -34.51 -8.30 7.15 -4.52 -10.04
Baichuan-13B-Chat 0.64 0.69 12.79 8.00 0.13 5.40
GPT-3.5-turbo-0613 0.65 -29.81 16.71 26.91 -39.69 -6.47
LLaMA-2-13B 0.57 29.19 38.00 34.54 - 33.91
LLaMA-2-7B 0.54 25.57 41.09 27.09 - 31.25
LLaMA-2-7B-Chat 0.61 25.39 23.81 12.52 - 20.58
Zhongjing-Base 0.59 25.16 37.70 14.65 - 25.83
InternLM-Chat-7B 0.48 3.43 25.13 21.65 - 16.74
Baichuan2-7B-Base 0.60 14.99 32.31 4.81 - 17.37
Baichuan2-7B-Chat 0.60 16.12 22.67 -18.59 - 6.73
Mistral-7B-v0.1 0.61 30.25 34.87 -22.79 - 14.11
Phi-1.5 0.65 19.30 10.78 -17.57 - 4.17
Baichuan2-13B-Base 0.61 10.81 17.67 6.75 - 11.75
Colossal-LLaMA-2-7B-Base 0.58 15.99 3.33 - - 9.66
Qwen-14B-Chat 0.59 14.68 14.30 - - 14.49
Qwen-7B 0.62 7.45 -7.31 - - 0.07
Qwen-7B-Chat 0.48 60.26 74.85 - - 67.55
Skywork-13B-Base 0.61 20.20 5.04 - - 12.62
ChatGLM3-6B 0.60 3.24 6.29 - - 4.76
Zephyr-7B-beta 0.61 14.33 18.22 - - 16.28
Yi-6B 0.63 8.42 -9.47 - - -0.53
Yi-6B-Chat 0.67 13.11 -6.19 - - 3.46
Qwen-1.8B 0.68 -6.93 -33.98 - - -20.46
Qwen-1.8B-Chat 0.73 4.74 -8.91 - - -2.08
RWKV-v5-Eagle-7B 0.70 -0.09 -20.38 - - -10.23
GPT-4-1106 0.42 6.67 -7.11 - - -0.22
TinyLLaMA-1.1B-Chat-v0.6 0.71 13.27 -16.30 - - -1.51
Gemini-pro-1213 0.71 -15.15 -41.28 - - -28.21
Phi-2 0.81 7.11 - - - 7.11
Table 9: Percentage Change in Model Accuracy Following Release
4.4 Correlation with Existing Benchmarks
Following our previous chapter 3.4, we continue to explore the correlation between prediction performance and existing benchmarks, using the same benchmarks as in 3.4. We have categorized the tags of questions into several groups: Societal Issues Items, Political and Government Affairs, Economic and Business Items, and Science and Technology Items. Each group represents a significant area of interest.
Notably, our analysis revealed that while the Economic and Business Items and Science and Technology Items on base models displayed relatively high initial correlations, there appears to be little to no relationship in Societal issues Items and Political and Government Affairs, we think such topics are influenced by a multitude of unpredictable factors such as social changes, policy adjustments, and public sentiment. These elements make accurate predictions particularly challenging, leading to the observed low correlation with performance metrics. Detailed figures are in Appendix H.
For a detailed breakdown of the categories and the number of entries in each, please refer to the Appendix G. This provides a clearer picture of the dataset’s composition and the scope of our analysis.
4.5 Community Call to Action
We urge the community to take note of the above phenomenon: models tend to have a better comprehension of earlier events. This observation supports the findings in  Cheng et al. (2024), which suggest that models are most familiar with knowledge up to 2020. This observation is contrary to our initial hypothesis that models may overfit current events, which is significantly more concerning. It suggests that caution is needed, especially for those hoping to use model judgments directly.
5 Conclusion
We rigorously define and qualify temporal generalization and bias within LLMs, uncovering significant challenges in their ability to predict and adapt to future contexts. The experiments reveal that LLMs often struggle with temporal biases, affecting their reliability in real-world applications, particularly in dynamic scenarios requiring accurate forecasting.
The evaluation framework we provide allows for an in-depth assessment of these models’ performance over time. The findings underscore the urgent need for improvements in LLM training and updating processes to enhance their adaptability and mitigate biases.
Limitations
Despite the system’s innovative approach to evaluating LLMs using the latest textual data, it is important to acknowledge certain limitations. One significant limitation is the system’s applicability primarily to open LLMs that provide access to their logits. This dependency arises because our evaluation mechanism relies on analyzing these logits to compute loss metrics. Consequently, our system may not be directly applicable to non-open LLMs, where such detailed internal outputs are not accessible. This constraint highlights the need for adaptable evaluation strategies that can cater to a broader range of LLMs, including those with closed architectures.
Ethics Statement
  • Handling User-Uploaded Content: We respect the privacy and copyright of content uploaded by users. All data provided by users will be used solely for the purpose of model evaluation and not for any other purposes. Reasonable measures will be taken to protect this data from unauthorized access, use, or disclosure.
Appendix A Details of Data Collection
The primary data sources include:
Appendix B Data Overview of One Crawl
Table 10 presented below offers a comprehensive overview of the diverse sources and categories of texts that were analyzed in one data crawl. Classified into distinct groups such as Academic STEM, Academic Non-STEM, Internet QA, and Internet articles, the table delineates the number of texts extracted from each source along with their average length. This includes detailed counts from different academic archives like arXiv, various categories from Internet question and answer platforms like Reddit and Quora, and articles from well-known online platforms including Wiki and BBC.
Category Source Count Length
Academic STEM arXiv Mathematics 50 42850
arXiv Computer Science 50 43336
arXiv Physics 50 42580
arXiv Statistics 50 42619
arXiv Electrical Engineering and Systems Science 50 43027
Academic Non-STEM arXiv Economics 50 42592
arXiv Quantitative Finance 50 43969
arXiv Quantitative Biology 50 42974
Code GitHub 30 21849
Internet QA Reddit 583 643
Quora Health 20 1150
Quora Movies 69 1349
Quora Mathematics 70 1365
Quora Technology 80 1674
Internet article Wiki 220 4091
Wattpad 38 6615
BBC 45 4792
Yahoo 41 5373
Appendix C Detailed results of BPC generalization on Wiki and BBC dataset
OPT-13B 0.450 2.273 -2.087 1.265 0.839 0.572
OPT-2.7B 0.475 1.970 -2.228 1.192 1.143 0.519
LLaMA-7B 0.450 1.270 0.177 -7.776 3.266 -0.766
Pythia-12B 0.480 -0.233 -3.888 -2.612 -1.012 -1.936
Falcon-rw-1B 0.510 0.459 -7.814 1.201 -1.474 -1.907
Baichuan-13B-Chat 0.467 -3.645 -1.078 0.413 - -1.436
LLaMA-2-13B 0.416 -6.928 3.502 1.683 - -0.581
LLaMA-2-7B 0.434 -7.314 3.092 1.279 - -0.981
LLaMA-2-7B-Chat 0.505 -8.308 2.695 0.664 - -1.650
Zhongjing-Base 0.465 -8.488 2.662 -1.290 - -2.372
InternLM-Chat-7B 0.515 -2.476 -1.344 0.236 - -1.195
Baichuan2-7B-Base 0.451 -3.059 0.595 -4.649 - -2.371
Baichuan2-7B-Chat 0.517 -3.571 -0.038 -5.315 - -2.975
Mistral-7B-v0.1 0.398 1.110 5.347 -0.661 - 1.932
Phi-1.5 0.602 -1.772 1.131 -1.837 - -0.826
Baichuan2-13B-Base 0.439 -2.943 0.882 -4.390 - -2.150
Baichuan2-13B-Chat 0.495 -3.541 -0.995 0.232 - -1.435
Colossal-LLaMA-2-7B-Base 0.618 2.825 3.715 - - 3.270
Qwen-14B-Chat 0.445 2.682 4.847 - - 3.764
Qwen-7B 0.441 1.892 3.689 - - 2.791
Qwen-7B-Chat 0.466 1.809 3.606 - - 2.708
Skywork-13B-Base 0.429 7.163 3.261 - - 5.212
ChatGLM3-6B 0.832 5.567 1.684 - - 3.626
Zephyr-7B-beta 0.435 7.340 3.928 - - 5.634
Yi-6B 0.427 6.478 3.144 - - 4.811
Yi-6B-Chat 0.446 3.030 -2.798 - - 0.116
Qwen-1.8B 0.520 1.899 -2.082 - - -0.091
Qwen-1.8B-Chat 0.599 1.635 -2.033 - - -0.199
RWKV-v5-Eagle-7B 0.440 4.956 2.278 - - 3.617
TinyLLaMA-1.1B-Chat-v0.6 0.513 2.539 -3.678 - - -0.569
Phi-2 0.487 1.882 - - - 1.882
Table 11: Percentage Change in BPC Following Release, measured by BBC news data from 2020 to 2024.3. Dash (”-”) symbols in the table indicate missing data points due to the later release dates of some models, which means there is no available data for those specific time intervals.
OPT-13B 0.075 -0.623 0.105 -0.413 -0.818 -0.437
OPT-2.7B 0.082 -0.658 0.024 -0.528 -0.915 -0.519
LLaMA-7B 0.052 0.214 -0.425 0.379 -0.067 0.025
Pythia-12B 0.071 -0.578 -0.933 -0.859 -0.903 -0.818
Falcon-rw-1B 0.083 -1.272 -1.113 -1.131 -1.207 -1.181
Baichuan-13B-Chat 0.058 0.019 0.124 -0.125 - 0.006
LLaMA-2-13B 0.041 1.700 1.446 1.796 - 1.647
LLaMA-2-7B 0.050 0.774 0.349 0.517 - 0.547
LLaMA-2-7B-Chat 0.060 0.132 -0.186 -0.093 - -0.049
Zhongjing-Base 0.050 0.756 0.269 0.569 - 0.531
InternLM-Chat-7B 0.080 -0.482 -0.301 -0.250 - -0.344
Baichuan2-7B-Base 0.059 0.523 0.157 0.185 - 0.289
Baichuan2-7B-Chat 0.068 0.343 -0.035 -0.073 - 0.078
Mistral-7B-v0.1 0.053 0.381 -0.028 0.018 - 0.124
Phi-1.5 0.098 0.066 -0.125 -0.248 - -0.102
Baichuan2-13B-Base 0.053 1.019 0.797 0.900 - 0.905
Baichuan2-13B-Chat 0.062 0.299 0.628 0.512 - 0.480
Colossal-LLaMA-2-7B-Base 0.065 -0.042 -0.199 - - -0.121
Qwen-14B-Chat 0.049 1.301 1.939 - - 1.620
Qwen-7B 0.061 0.428 0.430 - - 0.429
Qwen-7B-Chat 0.065 0.302 0.226 - - 0.264
Skywork-13B-Base 0.050 0.075 0.276 - - 0.175
ChatGLM3-6B 0.105 -0.187 -0.245 - - -0.216
Zephyr-7B-beta 0.059 -0.049 -0.033 - - -0.041
Yi-6B 0.058 -0.071 0.022 - - -0.024
Yi-6B-Chat 0.060 -0.028 0.023 - - -0.003
Qwen-1.8B 0.080 -0.131 -0.251 - - -0.191
Qwen-1.8B-Chat 0.091 -0.331 -0.496 - - -0.414
RWKV-v5-Eagle-7B 0.065 0.149 0.220 - - 0.184
TinyLLaMA-1.1B-Chat-v0.6 0.070 -0.332 -0.473 - - -0.403
Phi-2 0.072 -0.037 - - - -0.037
Table 12: Percentage Change in Language Likelihood Following Release, measured by wiki data from 2020 to 2024.3. Dash (”-”) symbols in the table indicate missing data points due to the later release dates of some models, which means there is no available data for those specific time intervals.
Appendix D Analysis of Language Likelihood Values and Benchmark Correlation Across Text and Model Groups
(a) STEM: base models
(b) non-STEM: base models
(c) Non-STEM: chat models
(a) STEM: all chat models
(b) STEM: 7B level chat models
Figure 6: (a) presents the correlations by loss for all 15 chat models on STEM Academia texts. (b) presents the correlations by loss for all 9 7B level chat models on STEM Academia texts. We can conclude that the scale of models being observed does not significantly alter outcomes, suggesting that performance correlations across benchmarks are generally consistent regardless of model size.
Appendix E Function Fitting Strategies
@@
Table 13: Variance of Residuals for Different Models
E.0.1 Findings
@@
Table 14: Model Scores
Appendix F Details of Model Accuracy Peaks
Table15 captures the peak performance periods of various models on Freshbench. It lists specific models along with their respective dates when they achieved the best performance.
@@ Appendix F Details of Model Accuracy Peaks
Table 15: Detailed Record of Model Accuracy Peaks
Appendix G Question Tag Details
@@
Appendix H Prediction Correlation over Time

### v2

@@ Appendix E TBI of Models across the arXiv, BBC, and Wiki datasets
Appendix F Most common tags of the Gjo questions
Table 8: Frequency of Topics
Topic Count
Business 648
Non-US Politics 453
Security and Conflict 446
Society 425
US Politics 363
Technology 359
Finance 299
Elections and Referenda 288
Economic Indicators 273
Foreign Policy 271
US Policy 247
Health 245
Leader Entry/Exit 219
Economic Policy 160
Sports 148
Entertainment 148
Environment 80
Open 16
Table 9: Most common tags of the Gjo questions
Appendix G Correlation with Existing Benchmarks
Following our previous chapter I, we continue to explore the correlation between prediction performance and existing benchmarks, using the same benchmarks as in I. We have categorized the tags of questions into several groups: Societal Issues Items, Political and Government Affairs, Economic and Business Items, and Science and Technology Items. Each group represents a significant area of interest.
Notably, our analysis revealed that while the Economic and Business Items and Science and Technology Items on base models displayed relatively high initial correlations, there appears to be little to no relationship in Societal issues Items and Political and Government Affairs, we think such topics are influenced by a multitude of unpredictable factors such as social changes, policy adjustments, and public sentiment. These elements make accurate predictions particularly challenging, leading to the observed low correlation with performance metrics. Detailed figures are in Appendix N.
Despite relatively high initial correlations in 2 question groups, there is a general lack of correlation between model predictions and existing benchmarks (mean correlation coefficient around zero). In addition, significant deviations occur in nearly half of the cases from early stages to subsequent periods. This suggests that the models’ understanding and alignment with benchmarks can vary substantially over time. So temporal generalization provides novel insights and draws attention to an aspect that has been often ignored by previous benchmarks.
Appendix H Analysis of Language Likelihood Values and Benchmark Correlation Across Text and Model Groups
(a) STEM: base models
(b) non-STEM: base models
(c) Non-STEM: chat models
(a) STEM: base models
(b) non-STEM: base models
(c) Non-STEM: chat models
Figure 5: The heatmaps display the correlations by loss for a group of models across different text chunk lengths and different benchmark scores. Each row represents a specific truncation length, while the columns represent different benchmark score categories. The values in the cells indicate the corresponding correlation coefficients. (a) presents the correlations by loss for all 18 base models on STEM Academia texts. (b) presents the correlations by loss for all 18 base models on non-STEM Academia texts. (c) presents the correlations by loss for all 15 chat models on non-STEM Academia
(a) STEM: all chat models
(b) STEM: 7B level chat models
Figure 6: (a) presents the correlations by loss for all 15 chat models on STEM Academia texts. (b) presents the correlations by loss for all 9 7B level chat models on STEM Academia texts.
From Fig. 6 we can conclude that the scale of models being observed does not significantly alter outcomes, suggesting that performance correlations across benchmarks are generally consistent regardless of model size.
Appendix I Correlation with Existing Benchmarks
@@
Table 10: Overview of AI Benchmarks and Their Evaluation Focus
@@
HellaSwag Tests an AI’s common sense reasoning by requiring it to complete sentences or narratives that reflect everyday logic, highlighting the model’s ability to predict logical continuations. [29]
MMLU Evaluates an AI’s breadth and depth of understanding across a diverse range of subjects, from science to literature, showcasing its versatility and comprehensive grasp of human language. [30]
TruthfulQA Focuses on the model’s capacity to provide honest and accurate responses, particularly to questions where the truth may not be intuitive, testing its commitment to truthfulness and its proficiency in avoiding misinformation. [31]
Winogrande Assesses common sense reasoning and natural language understanding through sentence completion tasks, measuring an AI’s ability to apply common sense knowledge in language processing. [32]
GSM8K Challenges AI models with elementary-level math problems, testing their mathematical reasoning capabilities and their understanding and application of basic math concepts. [33]
@@
Tab. 11 shows the comparative analysis of base and tuned models. Our findings suggest that: 1) Base models exhibit a stronger correlation with established benchmarks, implying that these models, in their fundamental form, possess a robust capacity for modeling language. This observation advocates for conducting initial evaluations using base models to capture their intrinsic language modeling capabilities. 2) The exploration of tuned models reveals potential deeper patterns in model tuning, suggesting that specific tuning strategies may unlock further enhancements in model performance across different tasks and domains.
Table 11: Correlation with Different Benchmarks, Base vs. Chat Models
@@
Tab. 12 is about a long text understanding benchmark Longbench.
@@
Table 12: Correlation with Longbench values across different lengths
@@
As can be seen in Appendix  H, whether all models are observed or only those of a specific scale does not make a big difference. Comparisons across different model sizes show that the differences are not substantial, indicating consistency in performance correlations across benchmarks.
Further analysis from Tab. 13 within the MMLU benchmark sections reveals interesting patterns. While we assume models to be generally excel in domains closely aligned with their training data, this correlation is not strong. Specifically, STEM scores do not show strong correlation with STEM-related texts. Moreover, academic papers, even in Social Sciences and Humanities, exhibit stronger correlations than Internet data, which includes forums and news.
@@
Appendix J Function Fitting Strategies
@@
J.0.1 Findings
@@
Appendix K Details of Model Accuracy Peaks
Tab. 16 captures the peak performance periods of various models on Freshbench. It lists specific models along with their respective dates when they achieved the best performance.
Table 16: Detailed Record of Model Accuracy Peaks
@@
Appendix L Question Tag Details
@@
Appendix M Analysis of Best-Achieved Accuracy and Cutoff Time
Peak Period Model Peak Time of Accuracy Cutoff Time Interpolation
Post-2019 TinyLLaMA1.1B 2019-07 2023-11 52
2017 to 2018 GPT4-1106 2018-07 2023-06 59
LLaMA2-13B 2018-07 2023-02 55
Baichuan2-13B-Base 2018-01 2023-06 65
2015 to 2016 Baichuan2-7B-Chat 2016-01 2023-08 91
Colossal-LLaMA-2-7B-Base 2016-01 2023-09 92
LLaMA2-7B 2016-01 2023-02 85
Tab. 18 shows an analysis covering the period from 2015 to 2024, with the year 2019 serving as a pivotal midpoint, a distinct pattern emerges from the performance data of various language models: all models reached their peak accuracy before 2020. This observation points to a notable trend toward "Nostalgia", where models demonstrate optimal performance on data from the earlier part of the timeline analyzed. This tendency highlights the models’ alignment with the informational characteristics prevalent prior to the latter half of the observed period, suggesting their training data may disproportionately reflect earlier times.
Appendix N Prediction Correlation over Time